# VCC-INFUSE: Towards Accurate and Efficient Selection of Unlabeled Examples in Semi-supervised Learning

### v1

@@ Abstract
Despite the progress of Semi-supervised Learning (SSL), existing methods fail to utilize unlabeled data effectively and efficiently. Many pseudo-label-based methods select unlabeled examples based on inaccurate confidence scores from the classifier. Most prior work also uses all available unlabeled data without pruning, making it difficult to handle large amounts of unlabeled data. To address these issues, we propose two methods: Variational Confidence Calibration (VCC) and Influence-Function-based Unlabeled Sample Elimination (INFUSE). VCC is an universal plugin for SSL confidence calibration, using a variational autoencoder to select more accurate pseudo labels based on three types of consistency scores. INFUSE is a data pruning method that constructs a core dataset of unlabeled examples under SSL. Our methods are effective in multiple datasets and settings, reducing classification errors rates and saving training time. Together, VCC-INFUSE reduces the error rate of FlexMatch on the CIFAR-100 dataset by 1.08% while saving nearly half of the training time.
@@ 3.1 Ensemble Consistency
where $p(y|x,\theta)$ represents the probability distribution of the label $y$ of $x$ given the parameters $\theta$, and $p(\theta|D)$ represents the probability distribution of the model parameters $\theta$ trained on the dataset $D$. A single model may provide incorrect predictions for example $x$ due to randomness and noise, even if the confidence is high. Considering the entire parameter space, if all model parameters yield consistent predictions for $x$, the result is more convincing. In this case, the prediction can be viewed as an ensemble of predictions from multiple models.
@@ 3.4 Approximation of Calibrated Confidence
According to the results in Guo et al. (2017), calibration errors mainly occur in the middle range of confidences, while samples with extremely low or high confidences tend to have smaller calibration errors. Therefore, we approximately treat the lowest/highest confidence score in $q$ as well-calibrated and employs interpolation to calculate the calibrated confidence scores for other examples. To further eliminate the unfairness between different categories, the interpolation operation only considers examples with the same pseudo labels as the current example $u$.
$\begin{split}&q^{\prime}=\{e\mid e\in q,\arg\max\tilde{s}^{conf}_{e}=\arg\max\tilde{s}^{conf}_{u}\},\\ &max\_score=\max_{u^{\prime}\in q^{\prime}}\left({s}_{u^{\prime}}\right),min\_score=\min_{u^{\prime}\in q^{\prime}}\left({s}_{u^{\prime}}\right),\\ &max\_conf=\max_{u^{\prime}\in q^{\prime}}\left(\tilde{s}_{u^{\prime}}^{conf}\right),max\_conf=\max_{u^{\prime}\in q^{\prime}}\left(\tilde{s}_{u^{\prime}}^{conf}\right),\end{split}$ (11)
@@ Appendix A Optimizing VAE in VCC
$\begin{split}&\log p_{\theta}(r|c,x)\\ &=\int_{z}q_{\phi}(z|c,x)\log p_{\theta}(r|c,x)\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{p_{\theta}(z|r,c,x)}\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\log\left(\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{p_{\theta}(z|r,c,x)}\frac{q_{\phi}(z|c,x)}{q_{\phi}(z|c,x)}\right)\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)(\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{q_{\phi}(z|c,x)}\\ &\qquad\qquad\qquad\qquad\qquad+\log\frac{q_{\phi}(z|c,x)}{p_{\theta}(z|r,c,x)})\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{q_{\phi}(z|c,x)}\mathrm{d}z\\ &\qquad\qquad\qquad\qquad\qquad+D_{KL}(q_{\phi}(z|c,x)\|p_{\theta}(z|c,r,x))\\ &\geq\int_{z}q_{\phi}(z|c,x)\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{q_{\phi}(z|c,x)}\mathrm{d}z,\end{split}$ (22)
@@
$\begin{split}&D_{KL}(q_{\phi}(z|c,x)\|p_{\theta}(z|c,x))\\ &=\int_{z}q_{\phi}(z|c,x)(\log q_{\phi}(z|c,x)-\log p_{\theta}(z|c,x))\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)(\log\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(z-\mu)^{2}}{2\sigma^{2}}}-\log\frac{1}{\sqrt{2\pi}}e^{-\frac{z^{2}}{2}})\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\Big{(}-\frac{(z-\mu)^{2}}{2\sigma^{2}}+\log\frac{1}{\sqrt{2\pi\sigma^{2}}}+\frac{z^{2}}{2}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\log\frac{1}{\sqrt{2\pi}}\Big{)}\mathrm{d}z\end{split}$
$\begin{split}&=-\int_{z}q_{\phi}(z|c,x)\log{\sigma}\mathrm{d}z+\int_{z}q_{\phi}(z|c,x)\frac{z^{2}}{2}\mathrm{d}z\\ &\qquad\qquad\qquad\qquad\qquad\qquad\quad-\int_{z}q_{\phi}(z|c,x)\frac{(z-\mu)^{2}}{2\sigma^{2}}\mathrm{d}z\\ &=-\log{\sigma}+\mathbb{E}_{z\sim q_{\phi}}\left[\frac{z^{2}}{2}\right]-\mathbb{E}_{z\sim q_{\phi}}\left[\frac{(z-\mu)^{2}}{2\sigma^{2}}\right]\\ &=-\log{\sigma}+\frac{1}{2}\big{(}(\mathbb{E}_{z\sim q_{\phi}}[z])^{2}+Var(z)\big{)}-\frac{1}{2\sigma^{2}}Var(z)\\ &=-\log{\sigma}+\frac{\mu^{2}+\sigma^{2}}{2}-\frac{1}{2},\end{split}$ (24)

### v2

@@ Abstract
Despite the progress of Semi-supervised Learning (SSL), existing methods fail to utilize unlabeled data effectively and efficiently. Many pseudo-label-based methods select unlabeled examples based on inaccurate confidence scores from the classifier. Most prior work also uses all available unlabeled data without pruning, making it difficult to handle large amounts of unlabeled data. To address these issues, we propose two methods: Variational Confidence Calibration (VCC) and Influence-Function-based Unlabeled Sample Elimination (INFUSE). VCC is a universal plugin for SSL confidence calibration, using a variational autoencoder to select more accurate pseudo labels based on three types of consistency scores. INFUSE is a data pruning method that constructs a core dataset of unlabeled examples under SSL. Our methods are effective in multiple datasets and settings, reducing classification error rates and saving training time. Together, VCC-INFUSE reduces the error rate of FlexMatch on the CIFAR-100 dataset by 1.08% while saving nearly half of the training time.
@@ 3.1 Ensemble Consistency
where $p(y|x,\theta)$ represents the probability distribution of the label $y$ of $x$ given the parameters $\theta$, and $p(\theta|D)$ represents the probability distribution of the model parameters $\theta$ trained on the dataset $D$. A single model may provide incorrect predictions, for example, $x$ due to randomness and noise, even if the confidence is high. Considering the entire parameter space, if all model parameters yield consistent predictions for $x$, the result is more convincing. In this case, the prediction can be viewed as an ensemble of predictions from multiple models.
@@ 3.4 Approximation of Calibrated Confidence
According to the results in Guo et al. (2017), calibration errors mainly occur in the middle range of confidences, while samples with extremely low or high confidences tend to have smaller calibration errors. Therefore, we approximately treat the lowest/highest confidence score in $q$ as well-calibrated and employ interpolation to calculate the calibrated confidence scores for other examples. To further eliminate the unfairness between different categories, the interpolation operation only considers examples with the same pseudo labels as the current example $u$.
$\begin{split}&q^{\prime}=\{e\mid e\in q,\arg\max\tilde{s}^{conf}_{e}=\arg\max\tilde{s}^{conf}_{u}\},\\ &max\_score=\max_{u^{\prime}\in q^{\prime}}\left({s}_{u^{\prime}}\right),min\_score=\min_{u^{\prime}\in q^{\prime}}\left({s}_{u^{\prime}}\right),\\ &max\_conf=\max_{u^{\prime}\in q^{\prime}}\left(\tilde{s}_{u^{\prime}}^{conf}\right),min\_conf=\min_{u^{\prime}\in q^{\prime}}\left(\tilde{s}_{u^{\prime}}^{conf}\right),\end{split}$ (11)
@@ Appendix A Optimizing VAE in VCC
$\begin{split}&\log p_{\theta}(r|c,x)\\ &=\int_{z}q_{\phi}(z|c,x)\log p_{\theta}(r|c,x)\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{p_{\theta}(z|r,c,x)}\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\log\left(\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{p_{\theta}(z|r,c,x)}\frac{q_{\phi}(z|c,x)}{q_{\phi}(z|c,x)}\right)\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)(\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{q_{\phi}(z|c,x)}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\log\frac{q_{\phi}(z|c,x)}{p_{\theta}(z|r,c,x)})\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{q_{\phi}(z|c,x)}\mathrm{d}z\\ &\qquad\qquad\qquad\qquad\qquad+D_{KL}(q_{\phi}(z|c,x)\|p_{\theta}(z|c,r,x))\\ &\geq\int_{z}q_{\phi}(z|c,x)\log\frac{p_{\theta}(r|c,z,x)p_{\theta}(z|c,x)}{q_{\phi}(z|c,x)}\mathrm{d}z,\end{split}$ (22)
@@
$\begin{split}&D_{KL}(q_{\phi}(z|c,x)\|p_{\theta}(z|c,x))\\ &=\int_{z}q_{\phi}(z|c,x)(\log q_{\phi}(z|c,x)-\log p_{\theta}(z|c,x))\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)(\log\frac{1}{\sqrt{2\pi\sigma^{2}}}e^{-\frac{(z-\mu)^{2}}{2\sigma^{2}}}-\log\frac{1}{\sqrt{2\pi}}e^{-\frac{z^{2}}{2}})\mathrm{d}z\\ &=\int_{z}q_{\phi}(z|c,x)\Big{(}-\frac{(z-\mu)^{2}}{2\sigma^{2}}+\log\frac{1}{\sqrt{2\pi\sigma^{2}}}+\frac{z^{2}}{2}\\ &\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-\log\frac{1}{\sqrt{2\pi}}\Big{)}\mathrm{d}z\end{split}$
$\begin{split}&=-\int_{z}q_{\phi}(z|c,x)\log{\sigma}\mathrm{d}z+\int_{z}q_{\phi}(z|c,x)\frac{z^{2}}{2}\mathrm{d}z\\ &\qquad\qquad\qquad\qquad\qquad\quad-\int_{z}q_{\phi}(z|c,x)\frac{(z-\mu)^{2}}{2\sigma^{2}}\mathrm{d}z\\ &=-\log{\sigma}+\mathbb{E}_{z\sim q_{\phi}}\left[\frac{z^{2}}{2}\right]-\mathbb{E}_{z\sim q_{\phi}}\left[\frac{(z-\mu)^{2}}{2\sigma^{2}}\right]\\ &=-\log{\sigma}+\frac{1}{2}\big{(}(\mathbb{E}_{z\sim q_{\phi}}[z])^{2}+Var(z)\big{)}-\frac{1}{2\sigma^{2}}Var(z)\\ &=-\log{\sigma}+\frac{\mu^{2}+\sigma^{2}}{2}-\frac{1}{2},\end{split}$ (24)